# FASTopoWM: Fast-Slow Lane Segment Topology Reasoning with Latent World Models

### v1

@@ Introduction
where denotes the first transformer layer, consisting of a self-attention module, a lane-attention module (Li et al. 2023b), and a feed-forward network. According to the classification confidence of , the lowest-ranked queries of are substituted with stream queries . The slow branch of the remaining transformer layer is then formulated as:
@@
In this manner, the historical queries and the newly initialized queries are supervised concurrently, preventing the model from becoming overly reliant on historical queries. The temporal propagation in the slow system benefits from well-trained initialized queries, particularly at the first frame. Meanwhile, the fast system’s initialized queries implicitly gain from temporal dynamics through shared decoder parameters with the slow system.
Training Loss
@@
where denotes the lane segment loss, which supervises the predicted lane segments using Hungarian matching (Li et al. 2023b). The overall loss function in FASTopoWM is defined as follows:
Experiments
Datasets and Metrics
@@
Datasets. We evaluate our method on the multi-view lane topology benchmark OpenLane-V2 (Wang et al. 2024a), which comprises two subsets. SubsetA is re-annotated from Argoverse2 (Wilson et al. 2023), while SubsetB is re-annotated from nuScenes (Caesar et al. 2020). Both subsets contain surround-view images collected from 1,000 scenes. SubsetA features seven camera views, whereas SubsetB includes only six. SubsetA provides annotations for both lane segments and their topology, while SubsetB contains annotations only for centerlines and topology. To generate boundary annotations for SubsetB, we apply a standardized lane width perpendicular to the centerline. For centerline perception evaluation on SubsetB, we approximate centerlines by averaging the coordinates of the left and right boundary lines. We re-train LaneSegNet, StreamMapNet, and SQD-MapNet with the same configuration to obtain their results on SubsetB.
@@
Metrics. We conduct evaluations on two tasks: lane segment perception on SubsetA and centerline perception on SubsetB. To assess lane quality, we adopt Chamfer Distance and Fréchet Distance under fixed thresholds of {1.0, 2.0, 3.0} meters. For lane segments, APls and APped are employed to evaluate the detection performance of road lanes and pedestrian crossings, respectively. The mean average precision (mAP) is calculated as the average of APls and APped. To evaluate topology reasoning, we report TOPlsls. Lane boundary classification accuracy is measured by Accb (Yang et al. 2025b). The evaluation protocol for centerline perception follows a similar approach as lane segments. Additionally, OLS (Wang et al. 2024a) is computed between DETl and TOPll.
@@
Implementation Details
The PV-to-BEV encoder is composed of a pre-trained ResNet-50 (He et al. 2016), an FPN (Lin et al. 2017), and BevFormer (Li et al. 2022). The BEV features has a resolution of 200100, covering a perception area of 50m 25m. The decoder follows the Deformable DETR architecture, where the standard cross-attention module is substituted with lane attention (Li et al. 2023b). It consists of 6 transformer layers. A total of 200 queries are used, with 30% reserved for temporal propagation in the slow system. Both the query and BEV world models employ 2 transformer layers each. To optimize memory usage on the GPU, we introduce average pooling within the BEV world model to reduce the input BEV resolution from 200100 to 10050. Then, we use bilinear interpolation to restore the output resolution to 200100. The centerline, left boundary, and right boundary are ordered sequence of 10 points. Training is performed over 24 epochs with a batch size of 8 on NVIDIA V100 GPUs. To stabilize the streaming process, the first 12 epochs are trained using single-frame inputs. The learning rate is initialized at and follows a cosine annealing schedule throughout training. We use the AdamW optimizer (Kingma and Ba 2015). The loss weights and are set to 1.0 and 0.3, respectively. During inference, the fast and slow systems can operate in parallel to generate predictions, or alternatively, inference can be performed using only one of the systems.
Main Results
@@
| Action | Tem. | |||
| mAP | APls | APped | ||
| None | 36.5 | 35.1 | 37.9 | 28.5 |
| Traj. | 31.2 | 30.8 | 31.5 | 24.5 |
| Pos. | 37.4 | 36.4 | 38.4 | 29.6 |
| Architecture | Tem. | |||
| mAP | APls | APped | ||
| Linear Projection | 35.2 | 34.2 | 36.2 | 28.5 |
| MLPs | 36.2 | 35.9 | 36.2 | 28.9 |
| Transformer | 37.4 | 36.4 | 38.4 | 29.6 |
Results on Centerline Perception. The results are presented in Tab. 2. Our fast system surpasses LaneSegNet by 3.1% in OLS. Our slow system outperforms TopoFormer by 4.7% in OLS. By introducing world models to capture temporal information, our slow system achieves a 4.5% improvement in OLS over the fast system.
Ablation Studies
The ablation studies are mainly conducted on SubsetA.
Ablation Study on Training Methods. As shown in Tab. 4, the baseline method shows a significant drop in single-frame detection performance. PrevPredMap (Peng et al. 2025) addresses this issue by randomly alternating between single-frame and temporal training modes. However, this alternating training reduces the effectiveness of temporal feature learning. In contrast, our fast-slow system enables parallel training of both modes, achieving improvements of 2.1% mAP and 1.1% TOPlsls in temporal detection.
Ablation Study on World Models. As shown in Tab. 5, model performance is relatively insensitive to the number of query world model layers. While deeper models offer slight gains, they reduce inference efficiency. We use two layers as a trade-off. Tab. 6 examines the impact of different action conditions. Without action conditioning, the world model can still predict future states, but localization accuracy degrades. Inspired by end-to-end driving methods, we also condition on future trajectories, which provide future positions and could reduce reliance on pose estimation. However, this leads to clear performance drops, likely due to limited trajectory regression accuracy and the model’s lack of agent-awareness and dynamics inputs (e.g., speed, steering). Conditioning on relative ego pose yields the best results. Finally, Tab. 7 compares network architectures: linear layers and MLPs fail to capture temporal dependencies, while stacked transformers achieve the best performance.
Qualitative Results
Fig. 3 provides a qualitative result comparison between baseline method and our FASTopoWM. The baseline method produces more misaligned endpoints, which confuses topology reasoning. It also suffers from missed detections and hallucinated results. In contrast, our method yields good lane segment perception with accurate topology reasoning.
Conclusion

### v4

@@ 3.4 Unified Fast-Slow Decoder
where denotes the first transformer layer, consisting of a self-attention module, a lane-attention module (Li et al., 2023b), and a feed-forward network. According to the classification confidence of , the lowest-ranked queries of are substituted with stream queries . The slow branch of the remaining transformer layer is then formulated as:
@@
In this manner, the historical queries and the newly initialized queries are supervised concurrently, preventing the model from becoming overly reliant on historical queries. The temporal propagation in the slow system benefits from well-trained initialized queries, particularly at the first frame. Meanwhile, the fast system’s initialized queries implicitly gain from temporal dynamics through shared decoder parameters with the slow system. More details about prediction heads can be found in the appendix.
4 Training Loss
@@
where denotes the lane segment loss, which supervises the predicted lane segments using Hungarian matching (Li et al., 2023b). The details of can be found in the appendix. The overall loss function in FASTopoWM is defined as follows:
5 Experiments
| Method | Venue | Temporal | mAP | FPS | ||||
| MapTR (Liao et al., 2022) | ICLR23 | No | 27.0 | 25.9 | 28.1 | - | - | 14.5 |
| MapTRv2 (Liao et al., 2023) | IJCV24 | No | 28.5 | 26.6 | 30.4 | - | - | 13.6 |
| TopoNet (Li et al., 2023a) | Arxiv23 | No | 23.0 | 23.9 | 22.0 | - | - | 10.5 |
| LaneSegNet (Li et al., 2023b) | ICLR24 | No | 32.6 | 32.3 | 32.9 | 25.4 | 45.9 | 14.7 |
| TopoLogic (Fu et al., 2025a) | NIPS24 | No | 33.2 | 33.0 | 33.4 | 30.8 | - | - |
| Topo2Seq (Yang et al., 2025a) | AAAI25 | No | 33.6 | 33.7 | 33.5 | 26.9 | 48.1 | 14.7 |
[…]
| StreamMapNet (Yuan et al., 2024) | WACV24 | Yes | 20.3 | 22.1 | 18.6 | 13.2 | 33.2 | 14.1 |
| SQD-MapNet (Wang et al., 2024b) | ECCV24 | Yes | 26.0 | 27.1 | 24.9 | 16.6 | 39.4 | 14.1 |
[…]
5.1 Datasets and Metrics
@@
Datasets. We evaluate our method on the multi-view lane topology benchmark OpenLane-V2 (Wang et al., 2024a), which comprises two subsets. SubsetA is re-annotated from Argoverse2 (Wilson et al., 2023), while SubsetB is re-annotated from nuScenes (Caesar et al., 2020). Both subsets contain surround-view images collected from 1,000 scenes. SubsetA features seven camera views, whereas SubsetB includes only six. SubsetA provides annotations for both lane segments and their topology, while SubsetB contains annotations only for centerlines and topology. To generate boundary annotations for SubsetB, we apply a standardized lane width perpendicular to the centerline. For centerline perception evaluation on SubsetB, we approximate centerlines by averaging the coordinates of the left and right boundary lines. We re-train LaneSegNet, StreamMapNet, and SQD-MapNet with the same configuration to obtain their results on SubsetB.
| Method | Venue | Temporal | |||
| VectorMapNet (Liu et al., 2023) | ICML23 | No | - | 3.5 | - |
| STSU (Can et al., 2021) | ICCV21 | No | - | 8.2 | - |
| MapTR (Liao et al., 2022) | ICLR23 | No | - | 15.2 | - |
| TopoNet (Li et al., 2023a) | Arxiv23 | No | 25.1 | 24.3 | 6.7 |
| TopoMLP (Wu et al., 2023) | ICLR24 | No | 36.2 | 26.6 | 19.8 |
| LaneSegNet (Li et al., 2023b) | ICLR24 | No | 38.7 | 27.5 | 24.9 |
| TopoLogic (Fu et al., 2025a) | NIPS24 | No | 36.2 | 25.9 | 21.6 |
| TopoFormer⋆ (Lv et al., 2025) | CVPR25 | No | 41.5 | 34.8 | 23.2 |
[…]
| StreamMapNet (Yuan et al., 2024) | WACV24 | Yes | 26.7 | 18.9 | 11.9 |
| SQD-MapNet (Wang et al., 2024b) | ECCV24 | Yes | 29.1 | 21.9 | 13.3 |
[…]
@@
Metrics. We conduct evaluations on two tasks: lane segment perception on SubsetA and centerline perception on SubsetB. To assess lane quality, we adopt Chamfer Distance and Fréchet Distance under fixed thresholds of {1.0, 2.0, 3.0} meters. For lane segments, APls and APped are employed to evaluate the detection performance of road lanes and pedestrian crossings, respectively. The mean average precision (mAP) is calculated as the average of APls and APped. To evaluate topology reasoning, we report TOPlsls. Lane boundary classification accuracy is measured by Accb (Yang et al., 2025b). The evaluation protocol for centerline perception follows a similar approach as lane segments. Additionally, OLS (Wang et al., 2024a) is computed between DETl and TOPll.
5.2 Implementation Details
The PV-to-BEV encoder is composed of a pre-trained ResNet-50 (He et al., 2016), an FPN (Lin et al., 2017), and BevFormer (Li et al., 2022). The BEV features has a resolution of 200100, covering a perception area of 50m 25m. The decoder follows the Deformable DETR architecture, where the standard cross-attention module is substituted with lane attention (Li et al., 2023b). It consists of 6 transformer layers. A total of 200 queries are used, with 30% reserved for temporal propagation in the slow system. Both the query and BEV world models employ 2 transformer layers each. To optimize memory usage on the GPU, we introduce average pooling within the BEV world model to reduce the input BEV resolution from 200100 to 10050. Then, we use bilinear interpolation to restore the output resolution to 200100. The centerline, left boundary, and right boundary are ordered sequence of 10 points. Training is performed over 24 epochs with a batch size of 8 on NVIDIA V100 GPUs. To stabilize the streaming process, the first 12 epochs are trained using single-frame inputs. More details about stream-based training can be found in the appendix. The learning rate is initialized at and follows a cosine annealing schedule throughout training. We use the AdamW optimizer (Kingma & Ba, 2015). The loss weights and are set to 1.0 and 0.3, respectively. During inference, the fast and slow systems can operate in parallel to generate predictions, or alternatively, inference can be performed using only one of the systems.
5.3 Main Results
Results on Centerline Perception. The results are presented in Tab. 3. Our fast system surpasses LaneSegNet by 3.1% in OLS. Our slow system outperforms TopoFormer by 4.8% in OLS. By introducing world models to capture temporal information, our slow system achieves a 4.5% improvement in OLS over the fast system.
5.4 Ablation Studies
The ablation studies are mainly conducted on SubsetA.
@@ 5.4 Ablation Studies
Ablation Study on Training Methods. As shown in Tab. 2(a), the baseline method shows a significant drop in single-frame detection performance. PrevPredMap (Peng et al., 2025) addresses this issue by randomly alternating between single-frame and temporal training modes. However, this alternating training reduces the effectiveness of temporal feature learning. In contrast, our fast-slow system enables parallel training of both modes, achieving improvements of 2.1% mAP and 1.1% TOPlsls in temporal detection.
Ablation Study on World Models. Tab. 2(b) examines the impact of different action conditions. Without action conditioning, the world model can still predict future states, but localization accuracy degrades. Inspired by end-to-end driving methods, we also condition on future trajectories, which provide future positions and could reduce reliance on pose estimation. However, this leads to clear performance drops, likely due to limited trajectory regression accuracy and the model’s lack of agent-awareness and dynamics inputs (e.g., speed, steering). Conditioning on relative ego pose yields the best results. Tab. 2(c) compares network architectures: linear layers and MLPs fail to capture temporal dependencies, while stacked transformers achieve the best performance. As shown in Tab. 2(d), we fix the number of transformer layers in one world model while varying it in the other. The model performance is relatively insensitive to the number of query world model layers. While deeper models offer slight gains, they reduce inference efficiency. We use two layers as a trade-off.
@@
| Action | Tem. | |||
| mAP | APls | APped | ||
| None | 36.5 | 35.1 | 37.9 | 28.5 |
| Traj. | 31.2 | 30.8 | 31.5 | 24.5 |
| Pos. | 37.4 | 36.4 | 38.4 | 29.6 |
| Arch. | Tem. | |||
| mAP | APls | APped | ||
| Linear | 35.2 | 34.2 | 36.2 | 28.5 |
| MLPs | 36.2 | 35.9 | 36.2 | 28.9 |
| Transformer | 37.4 | 36.4 | 38.4 | 29.6 |
@@
5.5 Qualitative Results
Fig. 4 provides a qualitative result comparison between baseline method and our FASTopoWM under different road structures. The baseline method produces more misaligned endpoints, which confuses topology reasoning. It also suffers from missed detections and hallucinated results. In contrast, our method yields good lane segment perception with accurate topology reasoning. Fig. 6 visualizes the comparison of temporal detection results across 5 frames. The baseline method fails to maintain temporal topological consistency, resulting in false detections, missed detections, and hallucinated topologies. Our method demonstrates robust temporally consistent topology reasoning results.
6 Conclusion
@@
Appendix A Appendix
A.1 Action Latent
In the ablation study, we try to use both relative pose and trajectory as conditions for the world models. Ultimately, conditioning on relative poses achieve the best performance.
Related Pose. The relative pose is defined as the transformation matrix that maps the rotation matrix R and translation vector t from the LiDAR coordinate system of the previous frame to that of the current frame. It can be formulated as:
Trajectory. The trajectory represents the future positions of the ego vehicle. We extract positions over a 3-second horizon (equivalent to 6 frames) and transform them into the current BEV coordinate system to generate 6 waypoints (Li et al., 2024). In the final five frames of a scene, some future positions may be unavailable; in such cases, interpolation is applied to complete the trajectory. The trajectory query is randomly initialized and refined via cross-attention with the BEV features and lane segment queries. A MLP is used to predict the trajectory from trajectory query. Then, the loss for trajectory prediction can be formulated as:
We concatenate the predicted waypoints with the query and BEV features along the channel dimension, and use the world model to predict the state of the next frame.
A.2 Transformation Loss
Transformation losses are applied to supervise both the stream query and the query world model, with the goal of minimizing projection errors across frame transitions. We employ MLPs to predict lane segment coordinate, lane segment class, boundary class and BEV mask from stream queries :
where indicates the centerline reference points for stream queries. InSigmod refers to the inverse sigmoid function, while Denorm stands for denormalize. represents the lateral distance from the centerline to both the left and right lanes. , , and are coordinates of lane segment, classes of centerline, boundary types and semantic BEV mask. Then, the transformation losses are represented as:
where the values of , , , , , , and are 1.0, 0.01, 1.0, 1.0, 0.025, 1.0 and 3.0.
A.3 Prediction Heads
After the decoder, we employ MLPs to predict lane segment coordinate, lane segment class, boundary class and BEV mask from the updated queries :
where denotes centerline reference points and indicates positional embedding. denotes the adjacency matrix that encodes the topological associations. The confidence threshold for the adjacency matrix is set at 0.5
A.4 Lane Segment Loss
Lane segment loss is proposed by LaneSegNet (Li et al., 2023b):
where the hyperparameters are defined as: , , , , and . is the L1 loss computed between the predicted vectorized lanes and the ground truth lanes. consists of a Cross-Entropy loss and a Dice loss used to supervise the BEV semantic mask. The classification losses and are used for lane segment classification. is a focal loss applied to supervise the topological relationship prediction.
A.5 Streaming Training
We adopt the streaming training strategy for temporal fusion. For each training sequence, we randomly divide it into 2 splits at the start of each training epoch to foster more diverse data sequences. During inference, we use the entire sequences. Suppose a batch contains N samples, each from a different scene, read in chronological order. Temporal fusion is performed by determining whether the current data and the previously read data belong to the same scene. To facilitate temporal fusion, we introduce several memory modules, including stream query memory, stream BEV memory, and stream reference point memory, which store the predictions from the preceding frame.
A.6 DEMO
See the supplementary material vis.gif file for details. The visualization results demonstrate that our predictions maintain robust temporal consistency, reflected in the stable alignment of lane segment coordinates and topological structures as the ego vehicle moves.
A.7 Limitation and future work
World models demonstrate strong capability in predicting the next-timestep BEV and transforming query information for common driving scenarios, such as stationary states, straight-line driving, and gentle turns. However, their generalization may be inadequate for rare scenarios involving rapid changes in ego-vehicle pose, such as high-curvature turns. To address this, we plan to introduce noise to simulate diverse ego-vehicle pose variations, thereby enhancing the temporal transformation capability of world models in such corner cases. Furthermore, we intend to integrate Vision Language Models (VLMs) to aggregate world model outputs for improved detection, and leverage linguistic descriptions to enhance lane topology reasoning and interpretability. Finally, we aim to establish rules based on lane topology to provide actionable safety recommendations for autonomous driving.
A.8 Use of LLM
In this paper, Large Language Model is used only for writing enhancement purposes.